1.On January 1, 2020, Bramble Corp. purchased a new machine for $4310000. The new machine has an estimated useful life of nine years and the salvage value was estimated to be $143000. Depreciation was computed using the sum-of-the-years'-digits method. What amount should be shown in Bramble's balance sheet at December 31, 2021, net of accumulated depreciation, for this machine?

a.$3476600

b.$4167000

c.$2735800

d.$2643200

2. Sheffield Corp. purchased a depreciable asset for $702000 on April 1, 2018. The estimated salvage value is $72000, and the estimated useful life is 5 years. The straight-line method is used for depreciation. What is the balance in accumulated depreciation on May 1, 2021 when the asset is sold?

a

$294000

b

$252000

c

$346500

d

$388500

3. Sunland Company purchased a truck at the beginning of 2020 for $110200. The truck is estimated to have a salvage value of $3700 and a useful life of 124000 miles. It was driven 22000 miles in 2020 and 30000 miles in 2021. What is the depreciation expense for 2021? (Round intermediate calculations to 3 decimal places, e.g. 0.245.)

a

$27556

b

$7100

c

$25770

d

$44668
Answer #1

1) Sum of year digit = 9+8+7+6+5+4+3+2+1 = 45

Accumulated depreciation on December 31,2021 = (4310000-143000)*17/45 = 1574200

Asset net of Accumulated depreciation = 4310000-1574200 = 2735800

So answer is c) $2735800

2) Accumulated depreciation = (702000-72000/5*3) = 378000

3) Depreciation expense = (110200-3700/124000) = 0.859*30000 = 25770

So answer is c) $25770